### v1

# Distal Causal Excursion Effects: Modeling Long-Term Effects of Time-Varying Treatments in Micro-Randomized Trials

Abstract
Micro-randomized trials (MRTs) play a crucial role in optimizing digital interventions. In an MRT, each participant is sequentially randomized among treatment options hundreds of times. While the interventions tested in MRTs target short-term behavioral responses (proximal outcomes), their ultimate goal is to drive long-term behavior change (distal outcomes). However, existing causal inference methods, such as the causal excursion effect, are limited to proximal outcomes, making it challenging to quantify the long-term impact of interventions. To address this gap, we introduce the distal causal excursion effect (DCEE), a novel estimand that quantifies the long-term effect of time-varying treatments. The DCEE contrasts distal outcomes under two excursion policies while marginalizing over most treatment assignments, enabling a parsimonious and interpretable causal model even with a large number of decision points. We propose two estimators for the DCEE—one with cross-fitting and one without—both robust to misspecification of the outcome model. We establish their asymptotic properties and validate their performance through simulations. We apply our method to the HeartSteps MRT to assess the impact of activity prompts on long-term habit formation. Our findings suggest that prompts delivered earlier in the study have a stronger long-term effect than those delivered later, underscoring the importance of intervention timing in behavior change. This work provides the critically needed toolkit for scientists working on digital interventions to assess long-term causal effects using MRT data.
Keywords: causal inference, distal causal excursion effect, distal outcome, estimating equations, longitudinal data analysis, micro-randomized trials
1 Introduction
Micro-randomized trials (MRTs) are widely used to develop and optimize digital adaptive interventions (Klasnja et al., 2015; Liao et al., 2016), with applications in health, public policy, education, and information systems (Walton et al., 2018; NeCamp et al., 2020; Liu et al., 2023; Breitwieser et al., 2024; Pieper et al., 2024). In an MRT, each participant is sequentially randomized among treatment options (such as receiving or not receiving a push notification encouraging behavioral change), often hundreds or even thousands of times per participant. These interventions directly target proximal outcomes (i.e., short-term behavioral responses to treatment, such as step count in the next 30 minutes), but ultimately aim to influence distal outcomes (i.e., long-term behaviors and habit formation).
To analyze MRT data, the causal excursion effect (CEE) was introduced to model the effect of time-varying treatments on proximal outcomes (Boruvka et al., 2018; Qian et al., 2021). The CEE, along with the associated estimation methods proposed in those works, has become the standard approach to conduct primary and secondary analyses for MRTs (Klasnja et al., 2018; Nahum-Shani et al., 2021; Bell et al., 2023; Wang et al., 2023; Arévalo Avalos et al., 2024). However, these approaches do not apply to distal outcomes, leaving key scientific questions unanswered. For example, in a physical activity study (Klasnja et al., 2015), do activity prompts that target step count in the short term lead to sustained habit formation? In a smoking cessation study (Battalio et al., 2021), do mindfulness reminders help participants internalize skills to manage cravings after the intervention ends? Current practice relies on behavioral theory to assume the existence of long-term effects; thus, statistical methods that can infer the long-term effects from MRT data are critically needed.
To address this gap, we propose the distal causal excursion effect (DCEE) to quantify the impact of time-varying treatments on a distal outcome measured at the end of an MRT. The DCEE is a contrast between the distal outcomes under two treatment policies that represent excursions (i.e., deviations) from the MRT treatment policy. The DCEE definition implicitly averages over the stochastically assigned treatment at most decision points, thus enabling a parsimonious causal effect model even with a large number of decision points; a related marginalization idea was used in defining CEE for proximal outcomes. This marginalization overcomes a fundamental challenge in MRT analysis: the treatments are randomized at hundreds or even thousands of decision points. Classical causal models, such as marginal structural models and structural nested mean models (Robins, 2000), perform well when the number of treatment occasions is limited. However, in MRTs with many decision points, these models become impractical unless restrictive structural assumptions are imposed—assumptions that are often unrealistic in the behavior change context (Rudolph et al., 2022; Zhang et al., 2022).
To estimate the DCEE, we develop two estimators—one with cross-fitting and one without—both robust to outcome model misspecification. We establish their asymptotic properties and validate their finite-sample performance through simulations. We apply the proposed method to the HeartSteps MRT to assess the long-term effect of activity suggestions on habit formation measured at the end of the study. Our analysis reveals that activity suggestions delivered earlier in the study, despite being farther from the distal outcome, have a larger effect compared to those delivered later. We discuss the implications of these findings for designing future digital interventions.
Section 2 gives the definition of the distal causal excursion effect. Section 3 presents the estimators and the asymptotic theory. Section 4 presents the simulation study. Section 5 presents the real data application. Section 6 concludes with discussion.
2 DCEE Definition
2.1 Notation: MRT with a Distal Outcome
Consider an MRT with $n$ individuals, each in the trial for $T$ decision points, where treatment is randomized. In the following, variables without the subscript $i$ correspond to a generic individual. Let $A_{t}$ denote the binary treatment assignment at decision point $t$, where $A_{t}=1$ indicates treatment and $A_{t}=0$ indicates no treatment. Let $X_{t}$ denote observations recorded between decision points $t-1$ and $t$. The distal outcome, denoted by $Y$, is measured at the end of the study and typically represents the primary health outcome of interest. The observed data trajectory for an individual is given by $O=(X_{1},A_{1},X_{2},A_{2},\ldots,X_{T},A_{T},Y)$. Overbars indicate sequences of variables; for example, $\bar{A}_{t}=(A_{1},A_{2},\ldots,A_{t})$. The history of observed information for an individual up to decision point $t$ (excluding $A_{t}$) is $H_{t}=(X_{1},A_{1},\ldots,X_{t-1},A_{t-1},X_{t})$. At each decision point $t$, $A_{t}$ is assigned according to a randomization probability $p_{t}(a|H_{t}):=P(A_{t}=a\mid H_{t})$ for $a\in\{0,1\}$. Data from individuals are assumed to be independent and identically distributed samples from the distribution $P_{0}$, and expectations are taken with respect to $P_{0}$ unless stated otherwise. The data structure is illustrated in Fig. 1.
[…]
Figure 1: MRT data structure with a distal outcome. To simplify notation, a thick arrow is used to denote arrows from all nodes on the left to all nodes on the right.
In designing the MRT, researchers may deem it unsafe or unethical to deliver push notifications at specific times, such as when a participant is driving. At such decision points, the participant is considered ineligible for randomization, and no treatment will be delivered. Formally, $X_{t}$ includes an indicator $I_{t}$, with $I_{t}=1$ denoting being eligible for randomization at decision point $t$, and $I_{t}=0$ otherwise. If $I_{t}=0$, then $A_{t}=0$ deterministically. In the MRT literature $I_{t}=0$ is also referred to as being unavailable for treatment (Boruvka et al., 2018). We adopt the terminology “ineligible” for clarity.
Let $\mathbb{P}_{n}$ denote the empirical mean over all individuals. For any positive integer $k$, define $[k]:=\{1,2,\ldots,k\}$. The superscript $\star$ is used to indicate quantities associated with the true data-generating distribution $P_{0}$. For a vector $\alpha$ and a vector-valued function $f(\alpha)$, the notation $\partial_{\alpha}f(\alpha):=\partial f(\alpha)/\partial\alpha^{T}$ denotes the Jacobian matrix, where the $(i,j)$-th entry corresponds to the partial derivative of the $i$-th entry of $f$ with respect to the $j$-th entry of $\alpha$.
2.2 Distal Causal Excursion Effect (DCEE)
To define the causal effect, we use the potential outcomes notation (Rubin, 1974; Robins, 1986). Lowercase letters represent instantiations (non-random values) of the corresponding capital letter random variable. For example, $a_{t}$ is an instantiation of treatment $A_{t}$. For every individual, denote by $X_{t}(\bar{a}_{t-1})$ the $X_{t}$ that would have been observed at decision point $t$ if the individual were assigned a treatment sequence of $\bar{a}_{t-1}$ prior to $t$. The potential outcome of $H_{t}$ under $\bar{a}_{t-1}$ is $H_{t}(\bar{a}_{t-1})=\{X_{1},a_{1},X_{2}(a_{1}),a_{2},\ldots,X_{t}(\bar{a}_{t-1})\}$. The potential outcome of $Y$ under $\bar{a}_{T}$ is $Y(\bar{a}_{T})$.
We use the term policy to refer to any decision rule (static or dynamic) of assigning $A_{t}$ given the history $H_{t}$ for $t\in[T]$ (Murphy, 2003). The MRT policy, denoted by $D:=(d_{1},\ldots,d_{T})$, is the rule in the MRT to stochastically assign $A_{t}$ for every $t\in[T]$. Specifically, $d_{t}(\cdot)$ is a stochastic mapping that maps $H_{t}$ to $\{0,1\}$, such that $d_{t}(H_{t})=a$ with probability $P(A_{t}=a\mid H_{t})$ for $a\in\{0,1\}$. For each $t$, we consider two alternative decision rules for assigning $A_{t}$ that will be contrasted in defining the causal effect: $d_{t}^{1}$ is the decision rule that always assigns $A_{t}=1$ unless $I_{t}=0$ (in which case $A_{t}=0$), and $d_{t}^{0}$ always assigns $A_{t}=0$. In other words, $d_{t}^{1}(H_{t})=I_{t}$, and $d_{t}^{0}(H_{t})=0$. For $a\in\{0,1\}$, let $D_{d_{t}=d_{t}^{a}}$ denote the policy obtained by replacing $d_{t}$ with $d_{t}^{a}$ in the MRT policy $D$:
$D_{d_{t}=d_{t}^{a}}:=(d_{1},\ldots,d_{t-1},d_{t}^{a},d_{t+1},\ldots,d_{T}).$ (1)
$D_{d_{t}=d_{t}^{1}}$ and $D_{d_{t}=d_{t}^{0}}$ are two excursions (Qian et al., 2021; Guo et al., 2021) from the MRT policy $D$ in that they deviate from $D$ at decision point $t$ to always assign $A_{t}=1$ (while respecting the eligibility constraint) or always assign $A_{t}=0$, respectively.
Let $S_{t}$ denote a subset of $H_{t}$, which are the moderators of interest based on the scientific question. For example, one can set $S_{t}=\emptyset$ for a fully marginal effect that averages over all moderators, $S_{t}=A_{t-1}$ for effect moderation by previous intervention, or $S_{t}=X_{t}$ for effect moderation by current covariates. Note that $S_{t}$ can be time-varying. We consider the following distal causal excursion effect (DCEE) of $A_{t}$ on $Y$:
$\tau(t,s):=\mathbb{E}\Big{\{}Y(D_{d_{t}=d_{t}^{1}})-Y(D_{d_{t}=d_{t}^{0}})\mid S_{t}(\bar{d}_{t-1})=s\Big{\}}.$ (2)
$\tau(t,s)$ captures the difference in the potential outcomes of $Y$ contrasting two excursion policies, $D_{d_{t}=d_{t}^{1}}$ and $D_{d_{t}=d_{t}^{0}}$, conditional on effect modifiers $S_{t}$. In other words, $\tau(t,s)$ contrasts what would have happened if a participant followed the MRT policy throughout the study except for decision point $t$ where they received $A_{t}=I_{t}$ versus $A_{t}=0$. Technically, the notation $\tau(t,s)$ should further depend on the choice of $S_{t}\subset H_{t}$, but we omit this notational dependence to avoid unnecessary complexity. The choice of $S_{t}$ will be clear from the context.
Below we present several examples with different data-generating distributions to illustrate the DCEE and highlight its scientific relevance for digital interventions. In all examples, we assume that a larger $Y$ is desirable, the error term $\epsilon$ is exogenous with mean 0, and we set $S_{t}=\emptyset$ to focus on fully marginal effects $\tau(t)$. A time-varying variable is exogenous if it is independent of its own history and the history of other variables. More general versions of these examples, along with detailed derivations of the DCEE, are provided in the Supplementary Material A. Importantly, the estimators proposed in Section 3 are valid as long as the assumptions in Theorem 2 hold; they do not require knowledge of most aspects of the data-generating process, such as how $Y$ depends on $A_{t}$ and $X_{t}$ or how $A_{t}$ influences $X_{t+1}$ and future eligibility $I_{t+1}$.
Example 1 (DCEE marginalizes over effect modifiers not in $S_{t}$).
Suppose that for all $t\in[T]$, the covariate $X_{t}$ is exogenous, all individuals are always eligible ($I_{t}\equiv 1$), and treatment $A_{t}\sim\text{Bern}(p)$ is exogenous. Suppose $Y=g(\bar{X}_{T})+\sum_{t=1}^{T}A_{t}(\alpha_{t}+\beta_{t}X_{t})+\epsilon$ for some unspecified $g$. Since we set $S_{t}=\emptyset$, the DCEE $\tau(t)$ represents the fully marginal effect of $A_{t}$, given by $\tau(t)=\alpha_{t}+\beta_{t}~{}\mathbb{E}(X_{t})$. This illustrates how the DCEE naturally marginalizes over effect modifiers not explicitly included in $S_{t}$.
Example 2 (DCEE captures user burden).
Consider the same data-generating process as Example 1, but instead suppose that there is a negative interaction between consecutive treatments: $Y=g(\bar{X}_{T})+\sum_{t=1}^{T}\beta_{t}A_{t}-\sum_{t=1}^{T-1}\alpha_{t}A_{t}A_{t+1}+\epsilon$ with $\alpha_{t}\geq 0$. Here, treatment effects diminish when treatments are delivered in succession, capturing user burden. The average-over-time effect is $\frac{1}{T}\sum_{t=1}^{T}\tau(t)=\frac{1}{T}\sum_{t=1}^{T}\{\beta_{t}-(\alpha_{t}+\alpha_{t-1})p\}$, where $\alpha_{0}:=0$. The term $-(\alpha_{t}+\alpha_{t-1})p$ reflects the deterioration of treatment effects due to user burden, which increases with the treatment probability $p$. Thus, the DCEE automatically accounts for user burden, and as a result different MRTs with different $p$ will yield different DCEE.
Example 3 (DCEE combines direct and indirect effects).
Consider a two-timepoint setting ($T=2$), where $X_{1}$ is exogenous, all individuals are always eligible ($I_{1}=I_{2}\equiv 1$), treatments $A_{1},A_{2}\sim\text{Bern}(p)$ are exogenous, and the treatment $A_{1}$ influences the time-varying covariate $X_{2}$ through $X_{2}\mid A_{1}\sim\text{Bern}(\rho_{0}+\rho_{1}A_{1})$. Suppose $Y=\gamma_{0}+\gamma_{1}X_{1}+\gamma_{2}X_{2}+\beta_{1}A_{1}+\beta_{2}A_{2}+\epsilon$. In this case, the DCEE at $t=1$ is $\tau(1)=\beta_{1}+\gamma_{2}\rho_{1}$. This expression captures both the direct effect of $A_{1}$ on $Y$ ($\beta_{1}$) and the indirect effect mediated through $X_{2}$ ($\gamma_{2}\rho_{1}$), demonstrating how the DCEE naturally combines mediation pathways in time-varying interventions.
Example 4 (DCEE captures treatment effects on future eligibility).
Consider a two-timepoint setting ($T=2$), where all individuals are eligible at $t=1$ ($I_{1}\equiv 1$), treatment at $t=1$ influences future eligibility: $I_{2}\mid A_{1}\sim\text{Bern}(\rho_{0}-\rho_{1}A_{1})$, and the treatment assignments follow $A_{1}\sim\text{Bern}(p)$ and $A_{2}\mid I_{2}=1\sim\text{Bern}(p)$. Suppose $Y=\beta_{0}+\beta_{1}A_{1}+\beta_{2}A_{2}+\epsilon$. Assume $\beta_{1},\beta_{2},\rho_{0},\rho_{1}>0$. In this case, the DCEE at $t=1$ is $\tau(1)=\beta_{1}-\beta_{2}p\rho_{1}$. Here, the term $-\beta_{2}p\rho_{1}$ captures the reduced effect of $A_{1}$ on $Y$ due to its negative impact on future eligibility. This demonstrates how the DCEE accounts for delayed effects of treatment via eligibility constraints, a key consideration in MRTs as eligibility may depend on past treatment history.
Remark 1 (Differences between DCEE and CEE).
The DCEE differs from the causal excursion effect (CEE) (Boruvka et al., 2018; Qian et al., 2021) in several key ways: (i) Outcome focus: DCEE focuses on the end-of-study distal outcome, whereas CEE focuses on short-term proximal outcomes. (ii) Policy comparison: DCEE contrasts treatment policies on the entire sequence $A_{1},\ldots,A_{T}$, while CEE contrasts policies up to time $t$ (or up to $A_{t+\Delta-1}$ for some prespecified $\Delta\geq 1$). (iii) Handling of eligibility: DCEE marginalizes over the eligibility indicator and contrasts policies that respect the eligibility constraints, whereas CEE conditions on a decision point being eligible. The distinction in eligibility handling (iii) is particularly important: DCEE accounts for the long-term consequences of treatment on future eligibility, meaning that a treatment that is highly effective but substantially reduces future eligibility may lead to an attenuated DCEE (see Example 4). This issue does not arise in CEE, as it focuses only on short-term proximal outcomes. ∎
Remark 2 (Differences between DCEE and other causal estimands).
The DCEE also differs from other causal estimands for time-varying treatments, such as those considered in marginal structural model (MSM) or the structural nested mean model (SNMM) (Robins, 2000). In an MRT setting, a MSM would model the expected $Y$ under fixed treatment policies $(a_{1},\ldots,a_{T})$, and a SNMM would model the effect of a treatment blip on $Y$ (i.e., contrasting $A_{t}=1$ versus 00) after removing all future treatments (i.e., after setting $A_{t+1}=\ldots=A_{T}=0$). Without strong structural assumptions such as the Markovian property which are likely infeasible in behavior change settings, both approaches will introduce more parameters than can be reliably estimated from MRT data. In contrast, the DCEE compares stochastic treatment policies that deviate from the MRT policy at only one decision point, allowing for parsimonious modeling while preserving interpretability. This idea is related to the longitudinal modified treatment policies framework (Díaz et al., 2023), but DCEE uniquely accounts for effect modification by time-varying covariates and eligibility constraints, which are central to digital intervention studies. ∎
2.3 Causal Assumptions and Identification
We make the following standard causal assumptions regarding data from an MRT.
Assumption 1.
  1. (i)
    (SUTVA.) There is no interference across individuals and the observed data equals the potential outcome under the observed treatment. As a result, $X_{t}=X_{t}(\bar{A}_{t-1})$ for $t\in[T]$ and $Y=Y(\bar{A}_{T})$.
  2. (ii)
    (Positivity.) There exists a positive constant $\tau>0$, such that if $P(H_{t}=h_{t},I_{t}=1)>0$ then $\tau<P(A_{t}=a\mid H_{t}=h_{t},I_{t}=1)<1-\tau$ for $a\in\{0,1\}$.
  3. (iii)
    (Sequential ignorability.) For $1\leq t\leq T$, the potential outcomes $\big{\{}X_{t+1}(\bar{a}_{t})$, $X_{t+2}(\bar{a}_{t+1})$, $\ldots,X_{T}(\bar{a}_{T-1}),Y(\bar{a}_{T}):\bar{a}_{T}\in\{0,1\}\times\{0,1\}\times\cdots\times\{0,1\}\big{\}}$ are conditionally independent of $A_{t}$ given $H_{t}$.
Positivity and sequential ignorability are guaranteed by the MRT design. SUTVA holds for most MRTs without the social aspect but will be violated if interference is present, that is, if the treatment assigned to one participant affects the potential outcome of another participant. In such settings, a framework that incorporates causal interference is needed (Hudgens & Halloran, 2008; Shi et al., 2022). We do not consider such settings here.
Let $p^{1}_{t}(a\mid H_{t})$ denote the probability mass function of $A_{t}$ conditional on $H_{t}$ under the decision rule $d_{t}^{1}$: $p^{1}_{t}(a\mid H_{t})=P_{A_{t}\sim d_{t}^{1}(H_{t})}(A_{t}=a\mid H_{t})$, i.e., $p^{1}_{t}(1\mid H_{t})=I_{t}$ and $p^{1}_{t}(0\mid H_{t})=1-I_{t}$. Similarly, let $p^{0}_{t}(a\mid H_{t})$ denote the distribution of $A_{t}$ under $d_{t}^{0}$, i.e., $p^{0}_{t}(1\mid H_{t})=0$ and $p^{0}_{t}(0\mid H_{t})=1$. In the Supplementary Material B, we prove the following identification result for $\tau(t,s)$.
Theorem 1.
Under Assumption 1, we have
$\tau(t,s)$ $=\mathbb{E}\bigg{\{}\frac{p_{t}^{1}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-\frac{p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y~{}\bigg{|}~{}S_{t}=s\bigg{\}}.$ (3)
$=\mathbb{E}\Big{[}\mathbb{E}\big{\{}Y\mid H_{t},A_{t}=d_{t}^{1}(H_{t})\big{\}}-\mathbb{E}\big{\{}Y\mid H_{t},A_{t}=d_{t}^{0}(H_{t})\big{\}}\mid S_{t}=s\Big{]}$ (4)
$=\mathbb{E}(I_{t})\mathbb{E}\Big{\{}\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=1)-\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=0)\mid S_{t}=s,I_{t}=1\Big{\}}.$ (5)
3 Robust Estimators for DCEE
Let $f(t,s)$ be a pre-specified $p$-dimensional feature vector that depends on the decision point index and the effect modifier value. We consider estimating the best linear projection of $\tau(t,s)$ on $f(t,s)$, averaged over all decision points. Specifically, the true parameter $\beta^{\star}\in\mathbb{R}^{p}$ is defined as
$\beta^{\star}=\arg\min_{\beta\in\mathbb{R}^{p}}\sum_{t=1}^{T}\omega(t)\mathbb{E}\Big{[}\{\tau(t,S_{t})-f(t,S_{t})^{T}\beta\}^{2}\Big{]},$ (6)
where $\omega(t)$ is a pre-specified weight function satisfying $\sum_{t=1}^{T}\omega(t)=1$. This is similar to linearly parameterizing $\tau(t,S_{t})=f(t,S_{t})^{T}\beta^{\star}$, but by defining $\beta^{\star}$ this way the estimator is interpretable even when the linear parameterization does not reflect the true data-generating distribution.
Researchers may set different $f(t,s)$ and different $\omega(t)$ based on the specific scientific question. For example, when $S_{t}=\emptyset$, setting $f(t,s)=(1,t)$ or $f(t,s)=(1,t,t^{2})$ models how the DCEE may vary linearly or quadratically over the decision point where the excursion takes place. One may also include basis functions of $t$ in $f(t,s)$ to allow flexibility (see Section 5 for such an example). When $S_{t}\neq\emptyset$, $f(t,s)$ can be either additive in $t,s$ or include their interaction. Regarding the weights $\omega(t)$, setting $\omega(t)=1/T$ for all $t$ ensures that all decision points contribute to the estimation of $\beta$ equally. However, if the goal is to estimate $\tau(t,s)$ at a particular $t=t_{0}$, one can set $\omega(t_{0})=1$ and $\omega(t)=0$ for all $t\neq t_{0}$.
Remark 3 ($\beta^{\star}$ as a weighted average).
When $\omega(t)=1/T$, $S_{t}=\emptyset$, and $f(t,s)=1$ for all $t\in[T]$, $\beta^{\star}$ can be shown to adopt the following form:
$\beta^{\star}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}(I_{t})\mathbb{E}\Big{\{}\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=1)-\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=0)\Big{\}}.$ (7)
This is a time-averaged, fully marginal (i.e., not conditional on any effect modifiers) excursion effect of $A_{t}$ on $Y$ when the participant is eligible, discounted by the probability of being eligible at each decision point. For general $S_{t}$ and $f(t,s)$, when $\omega(t)=1/T$, $\beta^{\star}$ can be interpreted as a weighted average of the moderated excursion effects:
$\beta^{\star}=$ $\bigg{[}\sum_{t=1}^{T}\mathbb{E}\big{\{}f(t,S_{t})f(t,S_{t})^{T}\big{\}}\bigg{]}^{-1}$
×$\times\sum_{t=1}^{T}\mathbb{E}(I_{t})\mathbb{E}\Big{[}\big{\{}\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=1)-\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=0)\big{\}}f(t,S_{t})\Big{]}.$ (8)
The weighted average form of $\beta^{\star}$ for more general settings is provided and proved in the Supplementary Material C, which immediately implies (7) and (8). ∎
The identification equation (3) motivates the following rudimentary estimating function for $\beta$:
$\xi(\beta):=\sum_{t=1}^{T}\omega(t)\bigg{\{}\frac{p_{t}^{1}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-\frac{p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-f(t,S_{t})^{T}\beta\bigg{\}}f(t,S_{t}).$ (9)
We further subtract from $\xi(\beta)$ its projection on the score functions of the treatment selection probabilities to obtain a more efficient estimating function (Robins, 1999):
$\phi(\beta,\mu):=\sum_{t=1}^{T}\omega(t)\bigg{[}I_{t}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}\Big{\{}Y-p_{t}(0\mid H_{t})\mu_{t}(H_{t},1)-p_{t}(1\mid H_{t})\mu_{t}(H_{t},0)\Big{\}}-f(t,S_{t})^{T}\beta\bigg{]}f(t,S_{t}).$ (10)
Here, $\mu=(\mu_{1},\ldots,\mu_{T})$ is an infinite-dimensional nuisance parameter with truth $\mu_{t}^{\star}(H_{t},A_{t})=\mathbb{E}(Y\mid H_{t},A_{t})$. A detailed derivation of (10) is in the Supplementary Material D.
The form of $\phi(\beta,\mu)$ motivates two estimators, $\hat{\beta}$ and $\tilde{\beta}$, as depicted in Algorithms 1 and 2. $\hat{\beta}$ does not use cross-fitting and $\tilde{\beta}$ uses cross-fitting (Chernozhukov et al., 2018).
Stage 1: Fit $\mathbb{E}(Y\mid H_{t},A_{t})$ for $t\in[T]$. Denote the fitted model by $\hat{\mu}_{t}(H_{t},A_{t})$. Denote $\hat{\mu}:=(\hat{\mu}_{1},\ldots,\hat{\mu}_{T})$.
Stage 2: Obtain $\hat{\beta}$ by solving $\mathbb{P}_{n}\phi(\beta,\hat{\mu})=0$ with $\phi$ defined in (10).
Algorithm 1 A two-stage estimator $\hat{\beta}$ (without cross-fitting)
Stage 1: Take a $K$-fold equally-sized random partition $(B_{k})_{k=1}^{K}$ of observation indices $[n]=\{1,\ldots,n\}$. Define $B_{k}^{c}=[n]\setminus B_{k}$ for $k\in[K]$. For each $k\in[K]$, use solely observations from $B_{k}^{c}$ and fit $\mathbb{E}(Y\mid H_{t},A_{t})$ for $t\in[T]$. The fitted models using $B_{k}^{c}$ are denoted by $\hat{\mu}_{kt}$, and let $\hat{\mu}_{k}:=(\hat{\mu}_{k1},\ldots,\hat{\mu}_{kT})$.
Stage 2: Obtain $\tilde{\beta}$ by solving $K^{-1}\sum_{k=1}^{K}\mathbb{P}_{n,k}\phi(\beta,\hat{\mu}_{k})=0$. Here $\mathbb{P}_{n,k}$ denotes empirical average over observations from $B_{k}$.
Algorithm 2 A two-stage estimator $\tilde{\beta}$ (with cross-fitting)
We establish the asymptotic normality for $\hat{\beta}$ and $\tilde{\beta}$ in Theorem 2, which is proved in the Supplementary Material E.
Theorem 2 (Asymptotic normality of $\hat{\beta}$ and $\tilde{\beta}$).
Suppose Assumption 1 hold and consider $\beta^{\star}$ defined in (6). Suppose $\hat{\mu}$ converges to some limit $\mu^{\prime}$ (not necessarily the true $\mu^{\star}$) in $L_{2}$. Under regularity conditions, we have
$\sqrt{n}(\hat{\beta}-\beta^{\star})\stackrel{{\scriptstyle d}}{{\to}}N(0,V)\text{ as }n\to\infty,$ (11)
with $V:=\mathbb{E}\{\partial_{\beta}\phi(\beta^{\star},\mu^{\prime})\}^{-1}\mathbb{E}[\{\phi(\beta^{\star},\mu^{\prime})\phi(\beta^{\star},\mu^{\prime})^{T}\}]\mathbb{E}\{\partial_{\beta}\phi(\beta^{\star},\mu^{\prime})\}^{-1,T}$. $V$ can be consistently estimated by $\mathbb{P}_{n}\{\partial_{\beta}\phi(\hat{\beta},\hat{\mu})\}^{-1}\mathbb{P}_{n}[\{\phi(\hat{\beta},\hat{\mu})\phi(\hat{\beta},\hat{\mu})^{T}\}]\mathbb{P}_{n}\{\partial_{\beta}\phi(\hat{\beta},\hat{\mu})\}^{-1,T}$.
For the estimator $\tilde{\beta}$ that uses cross-fitting, assume that $\hat{\mu}_{k}$ converges to some limit $\mu^{\prime}$ (not necessarily the true $\mu^{\star}$) in $L_{2}$ for each $k\in[K]$. Under regularity conditions, the asymptotic normality (11) holds with $\hat{\beta}$ replaced by $\tilde{\beta}$, in which case $V$ can be consistently estimated by
$\bigg{[}\frac{1}{K}\sum_{k=1}^{K}\mathbb{P}_{n,k}\big{\{}\partial_{\beta}\phi(\hat{\beta},\hat{\mu}_{k})\big{\}}\bigg{]}^{-1}\bigg{[}\frac{1}{K}\sum_{k=1}^{K}\mathbb{P}_{n,k}\big{\{}\phi(\hat{\beta},\hat{\mu}_{k})\phi(\hat{\beta},\hat{\mu}_{k})^{T}\big{\}}\bigg{]}\bigg{[}\frac{1}{K}\sum_{k=1}^{K}\mathbb{P}_{n,k}\big{\{}\partial_{\beta}\phi(\hat{\beta},\hat{\mu}_{k})\big{\}}\bigg{]}^{-1,T},$
where $\mathbb{P}_{n,k}$ denotes the empirical average over observations from $B_{k}$.
Remark 4 (Robustness).
$\hat{\beta}$ and $\tilde{\beta}$ are robust in the sense that their consistency and asymptotical normality hold with arbitrarily fitted nuisance parameter $\hat{\mu}$ that may not need to converge to the true $\mu^{\star}$. Furthermore, the asymptotic variance $V$ only involves $\mu^{\prime}$, the limit of $\hat{\mu}$, and otherwise does not depend on how $\hat{\mu}$ is fitted. This is because the estimating function $\phi(\beta,\mu)$ is globally robust (Cheng et al., 2023), which we prove in the Supplementary Material E. Of course, a better $\hat{\mu}$ will enhance the efficiency of $\hat{\beta}$ and $\tilde{\beta}$. Using nuisance parameters that are not required for identification to improve efficiency is a common technique in causal inference (e.g., Tsiatis et al., 2008; Lok, 2024; Van Lancker et al., 2024). ∎
4 Simulation
4.1 Generative Model and True Parameter Value
We construct a generative model that simultaneously incorporates the following common complications in actual MRTs: endogenous time-varying covariates that are influenced by past treatments and affect future treatment assignments, effect moderation by time-varying covariates, current treatments interacting with past treatments which reflects user burden, and eligibility constraints.
In the generative model, we set the total number of decision points per person to $T=30$. Data from different individuals are i.i.d. We describe below the data-generating process for a generic individual. $X_{t}$ is a time-varying continuous covariate that depends on $A_{t-1}$ and $X_{t-1}$: $X_{t}=\theta_{0}+\theta_{1}A_{t-1}+\theta_{2}X_{t-1}+\eta_{t}$ with $\eta_{t}\sim N(0,1)$ being an exogenous error. $Z_{t}$ is a time-varying binary covariate that depends on $A_{t-1}$ and $Z_{t-1}$: $Z_{t}\sim\text{Bernoulli}\{\text{expit}(\zeta_{0}+\zeta_{1}A_{t-1}+\zeta_{2}Z_{t-1})\}$. We set $X_{0}=Z_{0}=A_{0}=0$. The eligibility indicator is generated exogenously: $I_{t}\sim\text{Bernoulli}(0.8)$. The randomization probability for $A_{t}$ is $P(A_{t}=1\mid H_{t},I_{t}=1)=\text{expit}\{(t-T/2)/T+Z_{t}-0.5+X_{t}/6\}$, where the transformations of the variables make their influence on $A_{t}$ comparable. The outcome is generated by
$Y=\sum_{t=1}^{T}\xi_{t}\{g(X_{t}/12+0.5)+Z_{t}\}+\sum_{t=1}^{T}A_{t}(\alpha_{t}+\nu_{t}X_{t}+\gamma_{t}Z_{t}+\lambda_{t}A_{t-1})+\epsilon,$
where $g(\cdot)$ is nonlinear and equals the probability density function for $\text{Beta}(2,2)$ distribution. The observed data for a generic individual is $\{X_{t},Z_{t},I_{t},A_{t}:1\leq t\leq T\}\cup\{Y\}$, and the randomization probability $P(A_{t}=1\mid H_{t},I_{t}=1)$ is known as part of the trial design.
The parameter values are set as follows: $\theta_{0}=-0.5$, $\theta_{1}=\theta_{2}=0.5$; $\zeta_{0}=-1$, $\zeta_{1}=\zeta_{2}=1$; $\alpha_{t}=1+2(t-1)/(T-1)$; $\nu_{t}=1+(t-1)/(T-1)$; $\gamma_{t}=1+0.5(t-1)/(T-1)$; $\lambda_{t}=-1-(t-1)/(T-1)$; $\xi_{t}=1+(t-1)/(T-1)$. The $\lambda_{t}$ takes negative values, reflecting user burden (Example 2). The magnitudes of $\alpha_{t},\nu_{t},\gamma_{t},\lambda_{t},\xi_{t}$ all increase with $t$, characterizing a scenario where the variables more proximal to the distal outcome $Y$ have a larger impact on $Y$.
We consider two sets of estimands. We set $\omega(t)=1/T$. The first estimand corresponds to the fully marginal DCEE by setting $S_{t}=\emptyset$ in (2) and $f(t,S_{t})=1$ in (6), and the true parameter value is $\beta^{\star}_{1}=1.603$. The second set of estimands corresponds to the DCEE moderated by the binary covariate $Z_{t}$, i.e., through setting $S_{t}=Z_{t}$ and $f(t,S_{t})=(1,Z_{t})^{T}$. The true parameter values are $(\beta^{\star}_{2},\beta^{\star}_{3})=(1.207,0.881)$. The numerical computation of $\beta^{\star}_{1}$ and $(\beta^{\star}_{2},\beta^{\star}_{3})$ is detailed in the Supplementary Material F.
4.2 Proposed Estimator and Comparator Methods
We consider the proposed estimators $\hat{\beta}$ (the one without cross-fitting) and $\tilde{\beta}$ (with $K=5$ fold cross-fitting). In both implementations, the working models for the nuisance parameters $\mu_{t}(H_{t},1)$ and $\mu_{t}(H_{t},0)$ are pooled over $t\in[T]$ and are fitted using the generalized additive model (gam in R package mgcv). Each working model includes covariates $X_{t}$ and $Z_{t}$, where the continuous covariate $X_{t}$ is included as a penalized spline with cubic spline basis and the binary covariate $Z_{t}$ is included untransformed. These working models are misspecified because of omitted variables ($\mu_{t}$ depends on all the history variables in addition to $X_{t},Z_{t}$) and incorrect functional form (the dependence of $\mu_{t}$ on $X_{t}$ and $Z_{t}$ varies with $t$, which is not captured in the working model that pools over all $t\in[T]$). We introduce this misspecification to illustrate the robustness of the proposed estimator to model specificiation in $\hat{\mu}$ (Remark 4).
The DCEE is a novel estimand and no existing method in the literature is directly applicable. Nonetheless, we choose as comparators two approaches commonly used to analyze longitudinal data with time-varying treatments to illustrate the advantage of the proposed method in estimating the DCEE: the generalized estimating equations (GEE, Liang & Zeger, 1986) and the structural nested mean model (SNMM, Robins, 1994). To use GEE to estimate the marginal DCEE $\beta^{\star}_{1}$, we specify the mean model for $Y$ as a linear combination of $A_{t},X_{t},Z_{t}$, and we pool data over all $t\in[T]$ when fitting the model. The coefficient for $A_{t}$ (same across all $t$) is the GEE estimator for $\beta^{\star}_{1}$. To use GEE to estimate the moderated DCEE $\beta^{\star}_{2}$ and $\beta^{\star}_{3}$, we specify the mean model for $Y$ as a linear combination of $A_{t},X_{t},Z_{t},A_{t}\times Z_{t}$ and pool data over all $t\in[T]$ when fitting the model. The coefficients for $A_{t}$ and $A_{t}\times Z_{t}$ are the GEE estimators for $\beta^{\star}_{2}$ and $\beta^{\star}_{3}$, respectively. We use a working independence correlation matrix and the robust standard error estimator.
For SNMM, we used the g-estimation-based implementation in R package DTRreg. To estimate the marginal DCEE $\beta^{\star}_{1}$, we specify the blip model for each $t\in[T]$ to only include an intercept. The treatment assignment model is correctly specified with $X_{t}$ and $Z_{t}$ included in the model for $A_{t}$, and because DTRreg fits separate models for each $t\in[T]$, the dependence of $A_{t}$ on $t$ is correctly modeled as well. The treatment-free model for mean outcome after the treatment blips at and after $t$ are removed is specified as a linear model that depends on $X_{t}$ and $Z_{t}$. This choice of the treatment-free model mimics the nuisance model for $\mu_{t}$ in the proposed method, so as to make the SNMM modeling as comparable to the proposed method as possible. To use SNMM to estimate the moderated DCEE $\beta^{\star}_{2}$ and $\beta^{\star}_{3}$, we add $Z_{t}$ into the blip model.
4.3 Results
We conducted simulations for sample sizes $n=30,50,100,200,300,400,500$, each with 1000 replicates. Results are shown in Fig. 2. The proposed $\hat{\beta}$ and $\tilde{\beta}$ have satisfactory performance for sample size as small as 30: for all estimands, bias is close to 0 and the coverage of 95% confidence intervals is close to the nominal level. There is no substantial difference between the implementations with and without cross-fitting. The two comparator methods, GEE and SNMM, suffer from substantial bias and poor coverage, because they were not designed for estimating the DCEE.
Figure 2: The numerical performance of DCEE, DCEE-cf, GEE, and SNMM for the three estimands in terms of bias (top panel), standard deviation (middle panel), and the coverage of 95% confidence intervals.
5 Application
We use data from HeartSteps I, the first of a sequence of HeartSteps MRTs, which aimed to design mHealth interventions that support sedentary adults in reaching and sustaining recommended physical activity levels (Klasnja et al., 2015). In this application, we focus on the micro-randomized activity suggestions, which are one of the interventions tested in the MRT. $n=37$ participants were enrolled in the study for six weeks. Decision points occurred five times per day at pre-determined times, with a total of $T=210$ decision points per person. At each decision point, a participant was eligible to be randomized if they were not driving or walking and their phone had a stable internet connection. At every eligible decision point, a participant was randomized with probability 0.6 to receive an activity suggestion, which is a push notification suggesting brief walking or stretching activity, and with probability 0.4 to receive no intervention. Across all participants, approximately 80% of the decision points turned out to be eligible. Each participant wore a wristband tracker throughout the study, which continuously recorded their step count.
The long-term goal of these activity suggestions was to help participants develop a habit for being physically active. Therefore, in this illustration we consider as distal outcome the average daily step count in the last week of the study, a proxy for positive habit formation. We analyze the effect of the activity suggestions delivered in the first five weeks on this distal outcome measured in Week 6. We discuss the choice of this distal outcome at the end of this section.
Throughout this section we set $\omega(t)=1/T$. In fitting the nuisance parameters $\mu_{t}(H_{t},1)$ and $\mu_{t}(H_{t},0)$, we use the generalized additive model (gam in R package mgcv), pool data over all $t\in[T]$ when fitting the model, and include the step count in the 30 minutes prior to decision point $t$ and a location indicator for whether the person is at home or work (location indicator $=1$) or elsewhere (location indicator $=0$).
We first study the fully marginal excursion effect by setting $S_{t}=\emptyset$ and $f(t,s)=1$ in (6). The estimated effect is $\hat{\beta}=69$ (95% CI $=[-107,245]$). This indicates that when averaged over the intervention period under consideration (Weeks 1 through 5), sending an activity suggestion, compared to not sending one, does not have a detectable effect on the distal outcome measured in Week 6. We further investigate the effect moderation by the decision point index $t$, i.e., the length of time a participant has spent in the study, because prior studies have found a deteriorating effect over time on a short-term proximal outcome measured after each decision point (Klasnja et al., 2018). We set $S_{t}=\emptyset$, $f(t,s)=(1,t-1)^{T}$, and $\beta=(\beta_{0},\beta_{1})^{T}$. The transformation $t-1$ in $f(t,s)$ ensures the interpretation of $\beta_{0}$ as the effect on the distal outcome for the treatment at the first decision point. The result is $\hat{\beta}_{0}=344$ (95% CI $=[-74,761]$) and $\hat{\beta}_{1}=-3.4$ (95% CI $=[-7.5,0.7]$). This means that the activity suggestion at the beginning of the study possibly has a sizeable effect (increasing the average daily step count in Week 6 by 344 steps) but this effect likely deteriorates over time (due to a negative $\hat{\beta}_{1}$). This is an intriguing finding: The interventions delivered at the beginning of the study are more distant from the proxy for habit formation measured in Week 6, but their effects on habit formation are actually larger than the interventions sent later in the study.
To investigate whether the linearity assumption made in the effect moderation analysis [where we set $f(t,s)=(1,t-1)^{T}$] is reasonable, we conducted another moderation analysis where $f(t,s)$ contains an intercept and B-spline basis of 6 degrees of freedom. This allows the causal excursion effect of $A_{t}$ on $Y$ [i.e., $\tau(t,s)$] to vary flexibly with $t$. To enhance the interpretability of the result, we constructed the B-spline basis on $\text{day}_{t}:=\lfloor\frac{t-1}{5}\rfloor+1$. Fig. 3 shows that there is an increasing trend in the DCEE during the first week of the study. After the effect reached its peak by the end of Week 1, it starts to gradually decrease and crosses 0 at around the end of Week 2.
Figure 3: How the effect of a push notification on the distal outcome measured in Week 6 varies across different days of the study in the HeartSteps MRT. Solid curve represents the point estimate and shaded area represents 95% pointwise confidence intervals.
This finding suggests that early-in-the-study activity suggestions in the HeartSteps study may play a crucial role in shaping long-term behavior change, even if they are more distant in time to the distal outcome. One possible explanation is that interventions delivered early in the study help establish an initial momentum for behavior change, laying the foundation for participants to internalize and sustain new activity patterns over time. These early nudges may have a lasting impact by fostering psychological mechanisms such as self-efficacy, intrinsic motivation, or goal-setting behaviors.
In contrast, activity suggestions provided later in the study have diminishing returns, due to two possible explanations. First, some participants may have already developed a habit later in the study and thus rely less on external prompts. Second, some participants may have already disengaged due to message fatigue or perceived redundancy in the prompts. This aligns with theories of habit formation, which suggest that consistent reinforcement in the early stages is crucial for establishing automaticity, while interventions introduced too late may struggle to alter established behavioral patterns (Lally & Gardner, 2013).
This finding highlights the importance of timing in digital interventions. It suggests that future mHealth interventions should prioritize delivering more intensive or strategically timed support early in the intervention period to promote long-term habit formation.
Lastly, we note that the activity suggestions were also randomly delivered in Week 6, and thus the distal outcome was also influenced by these interventions. Our analysis is still valid in the presence of such dependence, but the distal outcome and the analysis result should be viewed in the context of the Week 6 interventions. In other words, had the micro-randomizations been stopped by the end of Week 5 and the participants continued to wear the tracker in Week 6 allowing the distal outcome to still be measured, the DCEE in that hypothetical setting could be different due to potential interactions between treatments in Week 6 and those in earlier weeks.
6 Discussion
We introduced the distal causal excursion effect (DCEE), a novel causal estimand for evaluating the long-term impact of time-varying treatments in MRTs. The DCEE enables a parsimonious and interpretable model for the causal effect even when treatments are randomized at hundreds or thousands of decision points. We developed two robust estimators, one with cross-fitting and one without, both of which are robust to misspecification of the outcome model. Through asymptotic theory, simulations, and an application to the HeartSteps MRT, we demonstrated the feasibility and practical utility of our approach. For HeartSteps, our findings reveal that early-in-the-study activity suggestions may have a stronger long-term impact on habit formation at the end of the study. This work provides the critically needed toolkit for scientists working on digital interventions to assess long-term causal effects using MRT data.
In DCEE, we set the reference policy to be the MRT policy, and the excursion policies are one-step excursions that only differ from the reference policy at decision point $t$. The DCEE definition can be straightforwardly generalized to other reference policies and excursion policies. However, the feasibility of estimation may not easily generalize depending on how different these policies are from the MRT policy under which the data is collected, due to the potential instability of inverse probability weights. For estimating such effects, methods to stabilize the inverse weights could be explored as a future direction.
Future work could extend our approach in several directions. First, our method assumes no unmeasured confounding and a correctly specified propensity score, and both assumptions are satisfied in MRTs but not necessarily in observational studies; incorporating doubly robust estimation and methods to address unmeasured confounding would improve applicability to observational studies. Second, the current linear model for the DCEE could be extended to data-adaptive models, especially if the set of moderators $S_{t}$ is large. Third, the current approach estimates an average causal effect across all participants; future research could extend the method to estimate heterogeneous person-specific treatment effects. Fourth, linking proximal and distal outcomes through mediation analysis could clarify the mechanisms driving long-term behavior change. Finally, future work could explore policy learning approaches to identify the optimal sequence of treatments that yield the greatest long-term benefit.
Supplementary Materials
Supplementary Material A contains the extensions and details for the DCEE examples in Section 2.2. Supplementary Material B proves the identification result Theorem 1. Supplementary Material C proves the weighted average form of $\beta^{\star}$ given in Remark 3. Supplementary Material D derives the estimating function $\phi(\beta,\mu)$ in Eq. 10. Supplementary Material E proves the asymptotic normality Theorem 2. Supplementary Material F details the numerical computation of true parameter values in simulation studies.
Appendix A Extensions and Details for the DCEE Examples in Section 2.2
We provide the detailed derivation for the four DCEE Examples in Section 2.2, where because Assumption 1 holds in all these examples, we can invoke Theorem 1 and in particular Eq. 5 to calculate $\tau(t)$. We also present more general versions of Example 3 and Example 4 to illustrate additional nuanced effects that are captured by the DCEE.
A.1 Example 1: DCEE marginalizes over effect modifiers not in $S_{t}$
Consider the data-generating process described in Example 1. Below we show that $\tau(t)=\alpha_{t}+\beta_{t}~{}\mathbb{E}(X_{t})$.
Given the exogeneity of $X_{t}$ and $A_{t}$, and that $I_{t}\equiv 1$, we have
$\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=1)-\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=0)=\alpha_{t}+\beta_{t}X_{t},$
and the form of $\tau(t)$ follows immediately from Eq. 5.
A.2 Example 2: DCEE captures user burden
Consider the data-generating process described in Example 2. Below we show that $\tau(s)=\beta_{s}-(\alpha_{s-1}+\alpha_{s})p$ with $\alpha_{0}:=0$, and the form of $\frac{1}{T}\sum_{t=1}^{T}\tau(t)$ in Example 2 follows immediately.
Given the exogeneity of $X_{t}$ and $A_{t}$, and that $I_{t}\equiv 1$, we have
$~{}~{}~{}~{}\mathbb{E}(Y\mid H_{s},A_{s})$
$=\mathbb{E}\{g(\bar{X}_{T})\mid H_{s},A_{s}\}+\sum_{t=1}^{T}\beta_{t}\mathbb{E}(A_{t}\mid H_{s},A_{s})-\sum_{t=1}^{T-1}\alpha_{t}\mathbb{E}(A_{t}A_{t+1}\mid H_{s},A_{s})$
$=\mathbb{E}\{g(\bar{X}_{T})\mid H_{s}\}+\sum_{t=1}^{s-1}\beta_{t}A_{t}+\beta_{s}A_{s}+\sum_{t=s+1}^{T}\beta_{t}\mathbb{E}(A_{t}\mid H_{s})$
$~{}~{}~{}~{}-\sum_{t=1}^{s-2}\alpha_{t}A_{t}A_{t+1}-\alpha_{s-1}A_{s-1}A_{s}-\alpha_{s}A_{s}\mathbb{E}(A_{s+1}\mid H_{s})-\sum_{t=s+1}^{T}\alpha_{t}\mathbb{E}(A_{t}A_{t+1}\mid H_{s}).$
Therefore,
$\mathbb{E}(Y\mid H_{s},A_{s}=1)-\mathbb{E}(Y\mid H_{s},A_{s}=0)=\beta_{s}-\alpha_{s-1}A_{s-1}-\alpha_{s}p.$
Plugging this into Eq. 5 and we have
$\tau(s)=\mathbb{E}(\beta_{s}-\alpha_{s-1}A_{s-1}-\alpha_{s}p)=\beta_{s}-(\alpha_{s-1}+\alpha_{s})p.$
Averaging over $s\in[T]$ and we have the form of $\frac{1}{T}\sum_{t=1}^{T}\tau(t)$ in Example 2.
A.3 Example 3 (extension): DCEE combines direct and indirect effects
We present a more general version of Example 3. Consider a two-timepoint setting ($T=2$), where $X_{1}$ is exogeneous, all individuals are always eligible ($I_{1}=I_{2}\equiv 1$), treatments $A_{1},A_{2}\sim\text{Bern}(p)$ are exogeneous, and the treatment $A_{1}$ influences the time-varying covariate $X_{2}$ through $X_{2}\mid A_{1}\sim\text{Bern}(\rho_{0}+\rho_{1}A_{1})$. Generalizing Example 3, we assume that the outcome is generated from $Y=\gamma_{0}+\gamma_{1}X_{1}+\gamma_{2}X_{2}+\beta_{1}A_{1}+\beta_{2}A_{2}+\alpha_{1}X_{1}A_{1}+\alpha_{2}X_{2}A_{2}+\epsilon$.
We derive $\tau(1)$ below. We have
$\mathbb{E}(Y\mid X_{1},A_{1})$ $=\gamma_{0}+\gamma_{1}X_{1}+\gamma_{2}\mathbb{E}(X_{2}\mid A_{1})+\beta_{1}A_{1}+\beta_{2}\mathbb{E}(A_{2})+\alpha_{1}X_{1}A_{1}+\alpha_{2}\mathbb{E}(X_{2}\mid A_{1})\mathbb{E}(A_{2})$
$=\gamma_{0}+\gamma_{1}X_{1}+(\gamma_{2}+\alpha_{2}p)(\rho_{0}+\rho_{1}A_{1})+\beta_{1}A_{1}+\beta_{2}p+\alpha_{1}X_{1}A_{1}.$
Therefore,
$\mathbb{E}(Y\mid X_{1},A_{1}=1)-\mathbb{E}(Y\mid X_{1},A_{1}=0)=(\gamma_{2}+\alpha_{2}p)\rho_{1}+\beta_{1}+\alpha_{1}X_{1}.$
Plugging this into Eq. 5 and we have
$\tau(1)=\beta_{1}+\alpha_{1}\mathbb{E}(X_{1})+(\gamma_{2}+\alpha_{2}p)\rho_{1}.$
We interpret each term in $\tau(1)$:
  • $\beta_{1}$: the direct of $A_{1}$ on $Y$ not moderated by $X_{1}$;
  • $\alpha_{1}\mathbb{E}(X_{1})$: the direct of $A_{1}$ on $Y$ moderated by $X_{1}$;
  • $\gamma_{2}\rho_{1}$: the indirect of $A_{1}$ on $Y$ mediated by $X_{2}$ that does not interact with $A_{2}$;
  • $\alpha_{2}p\rho_{1}$: the indirect of $A_{1}$ on $Y$ mediated by $X_{2}$ that interacts with $A_{2}$.
A.4 Example 4 (extension): DCEE captures treatment effects on future eligibility
We present a more general version of Example 4. Consider a two-timepoint setting ($T=2$), where all individuals are eligible at $t=1$ ($I_{1}\equiv 1$), treatment at $t=1$ influences future eligibility: $I_{2}\mid A_{1}\sim\text{Bern}(\rho_{0}-\rho_{1}A_{1})$, and the treatment assignments follow $A_{1}\sim\text{Bern}(p)$ and $A_{2}\mid I_{2}=1\sim\text{Bern}(p)$. For simplicity assume that there are no other covariates besides $I_{t}$. Generalizing Example 4, we assume that the outcome is generated from $Y=\beta_{0}+\beta_{1}A_{1}+\beta_{2}A_{2}-\alpha A_{1}A_{2}+\epsilon$. Assume $\beta_{1},\beta_{2},\alpha,\rho_{0},\rho_{1}>0$.
We derive $\tau(1)$ below. We have
$\mathbb{E}(Y\mid A_{1})$ $=\beta_{0}+\beta_{1}A_{t}+(\beta_{2}-\alpha A_{1})\mathbb{E}(A_{2}\mid A_{1}).$
For $\mathbb{E}(A_{2}\mid A_{1})$ we have
$\mathbb{E}(A_{2}\mid A_{1})=\mathbb{E}\{\mathbb{E}(A_{2}\mid I_{2},A_{1})\mid A_{1}\}=\mathbb{E}(I_{2}p\mid A_{1})=p(\rho_{0}-\rho_{1}A_{1}).$
Plugging into $\mathbb{E}(Y\mid A_{1})$ and we have
$\mathbb{E}(Y\mid A_{1})=\beta_{0}+\beta_{1}A_{t}+(\beta_{2}-\alpha A_{1})p(\rho_{0}-\rho_{1}A_{1}).$
Therefore,
$~{}~{}~{}~{}\mathbb{E}(Y\mid A_{1}=1)-\mathbb{E}(Y\mid A_{1}=0)$
$=\beta_{1}+(\beta_{2}-\alpha)p(\rho_{0}-\rho_{1})-\beta_{2}p\rho_{0}$
$=\beta_{1}-p\rho_{0}\alpha+p\rho_{1}\alpha-p\rho_{1}\beta_{2}.$
Plugging this into Eq. 5 and we have
$\tau(1)=\beta_{1}-p\rho_{0}\alpha+p\rho_{1}\alpha-p\rho_{1}\beta_{2}.$
We interpret each term in $\tau(1)$:
  • $\beta_{1}$: the direct of $A_{1}$ on $Y$;
  • $-p\rho_{0}\alpha$: the negative interaction between $A_{1}$ and $A_{2}$ not accounting for the influence of $A_{1}$ on $I_{2}$;
  • $p\rho_{1}\alpha$: the reduced negative interaction of $A_{1}\times A_{2}$ due to $A_{1}$ negatively impacting future eligibility $I_{2}$;
  • $-p\rho_{1}\beta_{2}$: the reduced effectiveness of $A_{2}$ on $Y$ due to $A_{1}$ negatively impacting future eligiblity $I_{2}$.
Appendix B Proof of Identification Result (Theorem 1)
We first state and prove a useful lemma.
Lemma 3.
We have
  • (a)
    $p_{t}^{a}\{d_{t}^{a}(H_{t})\mid H_{t}\}=1$.
  • (b)
    $p_{t}^{a}\{1-d_{t}^{a}(H_{t})\mid H_{t}\}=0$.
Proof of Lemma 3.
When $a=1$, by definition we have $p_{t}^{1}(1\mid H_{t})=I_{t}={\mathds{1}}(I_{t}=1)$ and $p_{t}^{1}(0\mid H_{t})=1-I_{t}={\mathds{1}}(I_{t}=0)$. This implies that $p_{t}^{1}(a\mid H_{t})={\mathds{1}}(I_{t}=a)$. We also have $d_{t}^{1}(H_{t})=I_{t}$. Therefore, $p_{t}^{1}\{d_{t}^{1}(H_{t})\mid H_{t}\}=p_{t}^{1}(I_{t}\mid H_{t})=1$.
When $a=0$, by definition we have $d_{t}^{0}(H_{t})=0$ and $p_{t}^{0}\{d_{t}^{0}(H_{t})\mid H_{t}\}=p_{t}^{0}(0\mid H_{t})=0$. This proves statement (a).
Statement (b) follows immediately from statement (a) and the fact that
$p_{t}^{a}\{d_{t}^{a}(H_{t})\mid H_{t}\}+p_{t}^{a}\{1-d_{t}^{a}(H_{t})\mid H_{t}\}=1.$
This completes the proof. ∎
We now prove Theorem 1.
Proof of Theorem 1.
Repeatedly using the law of iterated expectations, we have
$~{}~{}~{}\mathbb{E}\bigg{\{}\frac{p_{t}^{a}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y~{}\bigg{|}~{}S_{t}\bigg{\}}$ (12)
$=\mathbb{E}\bigg{[}\mathbb{E}\bigg{\{}\frac{p_{t}^{a}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y~{}\bigg{|}~{}H_{t}\bigg{\}}~{}\bigg{|}~{}S_{t}\bigg{]}$
[…]
$=\mathbb{E}[\mathbb{E}\{Y\mid H_{t},A_{t}=d_{t}^{a}(H_{t})\}\mid S_{t}]$ (13)
$=\mathbb{E}(\mathbb{E}[\mathbb{E}\{Y\mid H_{t},A_{t}=d_{t}^{a}(H_{t}),\bar{A}_{t+1:T}^{a_{t}=d_{t}^{a}(H_{t})}\}\mid H_{t},A_{t}=d_{t}^{a}(H_{t})]\mid S_{t})$
$=\mathbb{E}(\mathbb{E}[\mathbb{E}\{Y(\bar{d}_{t-1},d_{t}^{a},\bar{d}_{t+1:T})\mid H_{t},A_{t}=d_{t}^{a}(H_{t}),\bar{A}_{t+1:T}^{a_{t}=d_{t}^{a}(H_{t})}\}\mid H_{t},A_{t}=d_{t}^{a}(H_{t})]\mid S_{t})$ (14)
$=\mathbb{E}(\mathbb{E}[\mathbb{E}\{Y(\bar{d}_{t-1},d_{t}^{a},\bar{d}_{t+1:T})\mid H_{t},A_{t}=d_{t}^{a}(H_{t})\}\mid H_{t},A_{t}=d_{t}^{a}(H_{t})]\mid S_{t})$ (15)
$=\mathbb{E}[\mathbb{E}\{Y(\bar{d}_{t-1},d_{t}^{a},\bar{d}_{t+1:T})\mid H_{t},A_{t}=d_{t}^{a}(H_{t})\}\mid S_{t}]$
$=\mathbb{E}[\mathbb{E}\{Y(\bar{d}_{t-1},d_{t}^{a},\bar{d}_{t+1:T})\mid H_{t}\}\mid S_{t}]$ (16)
$=\mathbb{E}\{Y(\bar{d}_{t-1},d_{t}^{a},\bar{d}_{t+1:T})\mid S_{t}(\bar{d}_{t-1})\}.$ (17)
Here, Eq. 13 follows from Lemma 3, Eq. 14 follows from Assumption 1 (i), Eq. 15 follows from Assumption 1 (iii), Eq. 16 follows from Assumption 1 (iii), Eq. 17 follows from Assumption 1 (i).
The fact that Eqs. 12, 13 and 17 are equal to each other establishes Eqs. 3 and 4.
Finally, to prove Eq. 5, we have
$~{}~{}~{}\mathbb{E}[\mathbb{E}\{Y\mid H_{t},A_{t}=d_{t}^{1}(H_{t})\}-\mathbb{E}\{Y\mid H_{t},A_{t}=d_{t}^{0}(H_{t})\}\mid S_{t}]$
$=\mathbb{E}[\mathbb{E}\{Y\mid H_{t},A_{t}=d_{t}^{1}(H_{t})\}-\mathbb{E}\{Y\mid H_{t},A_{t}=d_{t}^{0}(H_{t})\}\mid S_{t},I_{t}=1]P(I_{t}=1)$
$~{}~{}~{}+\mathbb{E}[\mathbb{E}\{Y\mid H_{t},A_{t}=d_{t}^{1}(H_{t})\}-\mathbb{E}\{Y\mid H_{t},A_{t}=d_{t}^{0}(H_{t})\}\mid S_{t},I_{t}=0]P(I_{t}=0)$
$=\mathbb{E}\{\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=1)-\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=0)\mid S_{t},I_{t}=1\}P(I_{t}=1)+0,$
and this establishes Eq. 5.
The proof is thus completed. ∎
Appendix C The Weighted Average Form of $\beta^{\star}$
We state and prove a general theorem for the weighted average form of $\beta^{\star}$ defined in Eq. 6, which immediately implies Eqs. 7 and 8.
Theorem 4.
Under Assumption 1, for the $\beta^{\star}$ defined in Eq. 6 we have
$\beta^{\star}=$ $\bigg{[}\sum_{t=1}^{T}\omega(t)\mathbb{E}\{f(t,S_{t})f(t,S_{t})^{T}\}\bigg{]}^{-1}$
×$\times\sum_{t=1}^{T}\omega(t)\mathbb{E}(I_{t})\mathbb{E}\Big{[}\big{\{}\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=1)-\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=0)\mid S_{t},I_{t}=1\big{\}}f(t,S_{t})\Big{]}.$ (18)
Proof of Theorem 4.
Let $L(\beta)$ be the loss function in Eq. 6, i.e.,
$L(\beta):=\sum_{t=1}^{T}\omega(t)\mathbb{E}\Big{[}\{\tau(t,S_{t})-f(t,S_{t})^{T}\beta\}^{2}\Big{]}.$
The right hand side of Eq. 18 is the $\beta^{\star}$ that minimizes the quadratic form $L(\beta)$. We prove this now.
We have
$L(\beta)=\sum_{t=1}^{T}\omega(t)\mathbb{E}\{\tau(t,S_{t})^{t}-2\tau(t,S_{t})f(t,S_{t})^{T}\beta+\beta^{T}f(t,S_{t})f(t,S_{t})^{T}\beta\}.$
Taking the derivative with respect to $\beta$ and set to 0 at $\beta=\beta^{*}$, we have
$\frac{\partial L(\beta^{\star})}{\partial\beta^{T}}=\sum_{t=1}^{T}\omega(t)\mathbb{E}\{-2\tau(t,S_{t})f(t,S_{t})^{T}+2(\beta^{\star})^{T}f(t,S_{t})f(t,S_{t})^{T}\}=0.$ (19)
This implies that
$\beta^{\star}=\bigg{[}\sum_{t=1}^{T}\omega(t)\mathbb{E}\{f(t,S_{t})f(t,S_{t})^{T}\}\bigg{]}^{-1}\bigg{[}\sum_{t=1}^{T}\omega(t)\mathbb{E}\{\tau(t,S_{t})f(t,S_{t})\}\bigg{]}.$ (20)
Plugging Eq. 5 from Theorem 1 into Eq. 20 and we immediately get Eq. 18. This completes the proof. ∎
Appendix D Derivation of the Estimating Function $\phi(\beta,\mu)$ in Eq. 10
We restate necessary definition. As defined in Eq. 9,
$\xi(\beta):=\sum_{t=1}^{T}\omega(t)\bigg{\{}\frac{p_{t}^{1}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-\frac{p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-f(t,S_{t})^{T}\beta\bigg{\}}f(t,S_{t}).$ (21)
Motivated by Robins (1999), we subtract from $\xi(\beta)$ its projection on the score functions of the treatment selection probabilities to obtain a more efficient estimating function:
$\xi(\beta)-\sum_{u=1}^{T}\Big{[}\mathbb{E}\{\xi(\beta)\mid H_{u},A_{u}\}-\mathbb{E}\{\xi(\beta)\mid H_{u}\}\Big{]}.$ (22)
Define $\xi_{t}(\beta)$ to be a summand in Eq. 21 so that $\xi(\beta)=\sum_{t=1}^{T}\xi_{t}(\beta)$:
$\xi_{t}(\beta):=\omega(t)\bigg{\{}\frac{p_{t}^{1}(A_{t}\mid H_{t})-p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-f(t,S_{t})^{T}\beta\bigg{\}}f(t,S_{t}).$ (23)
Then Eq. 22 becomes
$~{}~{}~{}\sum_{t=1}^{T}\bigg{(}\xi_{t}(\beta)-\sum_{u=1}^{T}\Big{[}\mathbb{E}\{\xi_{t}(\beta)\mid H_{u},A_{u}\}-\mathbb{E}\{\xi_{t}(\beta)\mid H_{u}\}\Big{]}\bigg{)}$
$=\sum_{t=1}^{T}\bigg{(}\xi_{t}(\beta)-\Big{[}\mathbb{E}\{\xi_{t}(\beta)\mid H_{t},A_{t}\}-\mathbb{E}\{\xi_{t}(\beta)\mid H_{t}\}\Big{]}\bigg{)}$ (24)
$~{}~{}~{}-\sum_{t=1}^{T}\bigg{(}\sum_{1\leq u\leq T,u\neq t}\Big{[}\mathbb{E}\{\xi_{t}(\beta)\mid H_{u},A_{u}\}-\mathbb{E}\{\xi_{t}(\beta)\mid H_{u}\}\Big{]}\bigg{)}.$ (25)
Let $\phi_{t}(\beta,\mu_{t})$ be a summand in Eq. 10 so that $\phi(\beta,\mu)=\sum_{t=1}^{T}\phi_{t}(\beta,\mu_{t})$:
$\phi_{t}(\beta,\mu_{t}):=\omega(t)\bigg{[}I_{t}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}\Big{\{}Y-p_{t}(0\mid H_{t})\mu_{t}(H_{t},1)-p_{t}(1\mid H_{t})\mu_{t}(H_{t},0)\Big{\}}-f(t,S_{t})^{T}\beta\bigg{]}f(t,S_{t}).$ (26)
It follows from Lemma 6, which we will establish below, that Eq. 24 equals $\phi(\beta,\mu)=\sum_{t=1}^{T}\phi_{t}(\beta,\mu_{t})$, the proposed estimating function with improved efficiency. The terms in Eq. 25 cannot be analytically derived without imposing additional models on the relationship between current and lagged variables in the longitudinal trajectory, thus we omit them when deriving the improved estimating function. This same heuristic was employed in Cheng et al. (2023) and Bao et al. (2024). Therefore, it suffices to establish the following lemmas.
Lemma 5.
We have
  • (a)
    $p_{t}^{1}(A_{t}\mid H_{t})=A_{t}I_{t}+(1-I_{t})$.
  • (b)
    $p_{t}^{0}(A_{t}\mid H_{t})=1-A_{t}$.
  • (c)
    $p_{t}^{1}(A_{t}\mid H_{t})-p_{t}^{0}(A_{t}\mid H_{t})=A_{t}I_{t}+A_{t}-I_{t}$.
  • (d)
    $\dfrac{p_{t}^{1}(A_{t}\mid H_{t})-p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}=I_{t}\dfrac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}$.
Proof of Lemma 5.
When $I_{t}=1$, $p_{t}^{1}(1\mid H_{t})=1$ and $p_{t}^{1}(0\mid H_{t})=0$. When $I_{t}=0$, $p_{t}^{1}(1\mid H_{t})=0$ and $p_{t}^{1}(0\mid H_{t})=1$. Therefore, $p_{t}^{1}(A_{t}\mid H_{t})=A_{t}I_{t}+(1-I_{t})$. This proves (a).
For $p_{t}^{0}(a\mid H_{t})$, regardless of $I_{t}$, we have $p_{t}^{0}(1\mid H_{t})=0$ and $p_{t}^{0}(0\mid H_{t})=1$. Therefore, $p_{t}^{0}(A_{t}\mid H_{t})=1-A_{t}$. This proves (b).
(c) follows immediately from (a) and (b).
To prove (d), first note that $p_{t}(A_{t}\mid H_{t})=A_{t}p_{t}(1\mid H_{t})+(1-A_{t})\{1-p_{t}(1\mid H_{t})\}$. This combined with (b) gives
$\frac{p_{t}^{1}(A_{t}\mid H_{t})-p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}=\frac{A_{t}I_{t}+A_{t}-I_{t}}{A_{t}p_{t}(1\mid H_{t})+(1-A_{t})\{1-p_{t}(1\mid H_{t})\}},$
which takes values $\frac{1}{p_{t}(1\mid H_{t})}$, $-\frac{1}{1-p_{t}(1\mid H_{t})}$, and 0 when $I_{t}=A_{t}=1$, $I_{t}=1$ and $A_{t}=0$, and $I_{t}=0$ (in which case $A_{t}=0$), respectively. We can directly verify that $I_{t}\dfrac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}$ also takes these exact same values under these three scenarios, respectively. This proves (d).
This completes the proof. ∎
Lemma 6.
Fix $t\in[T]$. For $\xi_{t}(\beta)$ defined in Eq. 23 and $\phi_{t}(\beta,\mu_{t})$ defined in Eq. 26, we have
$\phi_{t}(\beta,\mu_{t}^{\star})=\xi_{t}(\beta)-\mathbb{E}\{\xi_{t}(\beta)\mid H_{t},A_{t}\}+\mathbb{E}\{\xi_{t}(\beta)\mid H_{t}\}.$ (27)
Proof of Lemma 6.
By definition of $\xi_{t}(\beta)$ and Lemma 5(d), we have
$\xi_{t}(\beta):=\omega(t)\bigg{\{}I_{t}\dfrac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}Y-f(t,S_{t})^{T}\beta\bigg{\}}f(t,S_{t}).$ (28)
Therefore,
$\mathbb{E}\{\xi_{t}(\beta)\mid H_{t},A_{t}\}=\omega(t)\bigg{\{}I_{t}\dfrac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}\mathbb{E}(Y\mid H_{t},A_{t})-f(t,S_{t})^{T}\beta\bigg{\}}f(t,S_{t}).$ (29)
We also have
$\mathbb{E}\{\xi_{t}(\beta)\mid H_{t}\}$ $=\mathbb{E}\{\xi_{t}(\beta)\mid H_{t},I_{t}=1\}I_{t}+\mathbb{E}\{\xi_{t}(\beta)\mid H_{t},I_{t}=0\}(1-I_{t})$
$=\mathbb{E}\{\xi_{t}(\beta)\mid H_{t},I_{t}=1,A_{t}=1\}P(A_{t}=1\mid H_{t},I_{t}=1)I_{t}$
$~{}~{}~{}+\mathbb{E}\{\xi_{t}(\beta)\mid H_{t},I_{t}=1,A_{t}=0\}P(A_{t}=0\mid H_{t},I_{t}=1)I_{t}$
$~{}~{}~{}+\mathbb{E}\{\xi_{t}(\beta)\mid H_{t},I_{t}=0,A_{t}=0\}(1-I_{t})$
$=\omega(t)\bigg{\{}I_{t}\frac{1}{p_{t}(1\mid H_{t},I_{t}=1)}~{}\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=1)~{}p_{t}(1\mid H_{t},I_{t}=1)$
$~{}~{}~{}+I_{t}\frac{-1}{p_{t}(0\mid H_{t},I_{t}=1)}~{}\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=0)~{}p_{t}(0\mid H_{t},I_{t}=1)$
$~{}~{}~{}+0-f(t,S_{t})^{T}\beta\bigg{\}}f(t,S_{t})$ (30)
$=\omega(t)\Big{[}I_{t}\{\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=1)-\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=0)\}-f(t,S_{t})^{T}\beta\Big{]}f(t,S_{t}),$ (31)
where Eq. 30 follows from the equivalent form of $\xi(\beta)$ in Eq. 28.
Putting together Eqs. 28, 29 and 31, we have
$~{}~{}~{}\xi_{t}(\beta)-\mathbb{E}\{\xi_{t}(\beta)\mid H_{t},A_{t}\}+\mathbb{E}\{\xi_{t}(\beta)\mid H_{t}\}$
$=\omega(t)\bigg{(}I_{t}\bigg{[}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}\{Y-\mathbb{E}(Y\mid H_{t},A_{t})\}+\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=1)$
$~{}~{}~{}-\mathbb{E}(Y\mid H_{t},I_{t}=1,A_{t}=0)\}\bigg{]}-f(t,S_{t})^{T}\beta\bigg{)}f(t,S_{t})$
$=\omega(t)\bigg{(}I_{t}\bigg{[}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}\{Y-\mu_{t}^{\star}(H_{t},A_{t})\}+\mu_{t}^{\star}(H_{t},1)-\mu_{t}^{\star}(H_{t},0)\bigg{]}-f(t,S_{t})^{T}\beta\bigg{)}f(t,S_{t})$ (32)
$=\omega(t)\bigg{[}I_{t}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}\Big{\{}Y-p_{t}(0\mid H_{t})\mu_{t}^{\star}(H_{t},1)-p_{t}(1\mid H_{t})\mu_{t}^{\star}(H_{t},0)\Big{\}}-f(t,S_{t})^{T}\beta\bigg{]}f(t,S_{t}),$ (33)
where Eq. 32 follows from the definition of $\mu_{t}^{\star}(H_{t},A_{t})$, and Eq. 33 follows from directly verifying the equality when $A_{t}=1$ and when $A_{t}=0$. Eq. 33 is $\phi_{t}(\beta,\mu_{t}^{\star})$ defined in Eq. 26. This completes the proof. ∎
Appendix E Proof of Asymptotic Normality (Theorem 2)
We first state and proof a useful lemma.
Lemma 7.
We have
$\mathbb{E}\bigg{\{}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}~{}\bigg{|}~{}H_{t},I_{t}=1\bigg{\}}=0.$
Proof of Lemma 7.
By direct calculation we have
$~{}~{}~{}\mathbb{E}\bigg{\{}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}~{}\bigg{|}~{}H_{t},I_{t}=1\bigg{\}}$
$=\mathbb{E}\bigg{\{}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}~{}\bigg{|}~{}H_{t},I_{t}=1,A_{t}=1\bigg{\}}p_{t}(1\mid H_{t})+\mathbb{E}\bigg{\{}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}~{}\bigg{|}~{}H_{t},I_{t}=1,A_{t}=0\bigg{\}}p_{t}(0\mid H_{t})$
$=\frac{1}{p_{t}(1\mid H_{t})}p_{t}(1\mid H_{t})+\frac{-1}{p_{t}(0\mid H_{t})}p_{t}(0\mid H_{t})=0.$
This completes the proof. ∎
Now we prove Theorem 2.
Proof of Theorem 2.
We show that the estimating function $\phi(\beta,\mu)$ is globally robust, i.e., $\mathbb{E}\{\phi(\beta^{\star},\mu)\}=0$ for all $\mu$. Then, the asymptotic normality of $\hat{\beta}$ and $\tilde{\beta}$ follows immediately from Theorems 5.1 and 5.2 of Cheng et al. (2023).
By the definition of $\phi$ in Eq. 10, we have
$~{}~{}~{}\mathbb{E}\{\phi(\beta^{\star},\mu)\}$
$=\sum_{t=1}^{T}\omega(t)\mathbb{E}\bigg{[}\bigg{\{}I_{t}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}Y-f(t,S_{t})^{T}\beta^{\star}\bigg{\}}f(t,S_{t})\bigg{]}$
$-\sum_{t=1}^{T}\omega(t)\mathbb{E}\bigg{(}\bigg{[}I_{t}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}\Big{\{}p_{t}(0\mid H_{t})\mu_{t}(H_{t},1)+p_{t}(1\mid H_{t})\mu_{t}(H_{t},0)\Big{\}}\bigg{]}f(t,S_{t})\bigg{)}.$ (34)
The second term in Eq. 34 is 0 for any $\mu_{t}$ due to Lemma 7 and the law of iterated expectations. For the first term in Eq. 34, we have
$~{}~{}~{}\mathbb{E}\bigg{\{}I_{t}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}Yf(t,S_{t})\bigg{\}}$
$=\mathbb{E}\bigg{\{}\frac{p_{t}^{1}(A_{t}\mid H_{t})-p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Yf(t,S_{t})\bigg{\}}$ (35)
$=\mathbb{E}\bigg{[}\mathbb{E}\bigg{\{}\frac{p_{t}^{1}(A_{t}\mid H_{t})-p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y~{}\bigg{|}~{}S_{t}\bigg{\}}f(t,S_{t})\bigg{]}$
$=\mathbb{E}\{\tau(t,S_{t})f(t,S_{t})\},$ (36)
where Eq. 35 follows from Lemma 5(d) and Eq. 36 follows from Theorem 1. Plugging Eq. 36 to the first term in Eq. 34, we have
$~{}~{}~{}\sum_{t=1}^{T}\omega(t)\mathbb{E}\bigg{[}\bigg{\{}I_{t}\frac{(-1)^{1-A_{t}}}{p_{t}(A_{t}\mid H_{t})}Y-f(t,S_{t})^{T}\beta^{\star}\bigg{\}}f(t,S_{t})\bigg{]}$
$=\sum_{t=1}^{T}\omega(t)\mathbb{E}\{\tau(t,S_{t})f(t,S_{t})-f(t,S_{t})^{T}\beta^{\star}f(t,S_{t})\}=0,$
where the last equality is Eq. 19. Therefore, we have proved that $\mathbb{E}\{\phi(\beta^{\star},\mu)\}=0$ for all $\mu$. We now invoke Theorems 5.1 and 5.2 of Cheng et al. (2023) to immediately get Theorem 2. ∎
Appendix F Numerical Computation of True Parameter Values in Simulation Studies
The computation uses Eq. 20. In particular, for each $t\in[T]$, we generated two data sets under two excursion policies, each of sample size 1 million: one excursion policy is $D_{d_{t}=d_{t}^{1}}$ and the other is $D_{d_{t}=d_{t}^{0}}$. $\tau(t,S_{t}=\emptyset)$ is computed as the difference in means of $Y$ from the two data sets. $\tau(t,Z_{t}=0)$ and $\tau(t,Z_{t}=1)$ are computed as the difference in means of $Y$ from the two data sets among individuals with $Z_{t}=0$ and with $Z_{t}=1$, respectively. Then, a third data set of sample size 1 million is generated under the MRT policy, and the expectations in Eq. 20 are computed separately for the two sets of estimands using this third data set. This yields the numeric values of $\beta^{\star}_{1}$ and $(\beta^{\star}_{2},\beta^{\star}_{3})$.